### v1

@@ Appendix B Proofs of Auxiliary Propositions
Proof of subsection 2.1.
First, we need to confirm that (2.10) admits a unique classical solution on $[0,T]$. To prove existence, let function
@@
where $A(t)$ is any primitive function for $\frac{2a(t)}{c}$, so (2.11) gives a unique solution on $[0,T]$. The solution to (2.12) obviously exists and is unique, once we integrate both sides from $T$ to $t$. The proof is complete. ∎
Proof of subsection 2.1.
@@
We introduce the following notation: the Frobenius norm of any matrix $L$ is denoted by $\|L\|_{\mathrm{F}}$. It is clear that $\|M(t)\|_{\mathrm{F}}$ has a uniform bound for $t\in[0,T]$; we denote this bound by $M^{*}$, and denote $\|\bar{k}(t)\|_{\mathrm{F}}$ by $n(t)$. Multiply both sides of (2.13) with $2\bar{k}(t)^{\top}$, it follows that
@@
By basic ODE theorem, $n(t)$ stays below a function $N(t)$ characterized by
@@
By a simple calculation, $N(t)$ has a uniform bound for $t\in[0,T]$, so $\bar{k}(t)$ stays finite on $[0,T]$. Therefore, the existence of the solution holds true. Moreover, $M(t)$ is Lipschitz continuous on $[0,T]$, so the uniqueness of the solutions naturally follows from the Picard-Lindelöf theorem. The proof is complete.
Proof of subsection 2.1.
We adopt the notation established in the proofs of subsection 2.1. Deducing in the same way as in the proof of subsection 2.1, we obtain the existence property. Recalling that $M(t)$ is Lipschitz continuous, we obtain the uniqueness property. ∎
Proof of subsection 2.1.
We adopt the notation established previously in subsection 2.1. By the definition of $\phi_{b_{0}}$, as shown in subsection 2.1 , we discover that for any $b_{0}\neq 0$, $\frac{\phi_{b_{0}}-\phi_{0}}{b_{0}}$ solves (2.15). Consequently, from the uniqueness property provided by subsection 2.1, we have
@@
Let $t=T$. From $B_{1}(T)\neq 0$, we conclude that there exists a unique $b_{0}\in\mathbb{R}$ that satisfies $\phi_{b_{0}}(T)=-\zeta\gamma$. Thus, there exists only one solution to (2.13) and (2.14). The proof is complete. ∎
Proof of subsection 2.2.
@@
Note that $G(t)$ is continuously differentiable for every $t\in[0,T]$, so the solution of each equation, if exists, must be unique. To prove existence, we observe that $\|G(t)\|_{\mathrm{F}}$ has a uniform bound for $t\in[0,T]$, and work exactly as in subsection 2.1 to finish the proof. ∎
Proof of subsection 2.2.
Recalling from subsection 2.2 that $G(t)$ is continuously differentiable everywhere, we conclude that the solution, if exists, must be unique, by the Picard-Lindelöf theorem. The existence holds by the same way as in subsection 2.1. ∎
Proof of subsection 2.2.
We observe that , for $\forall b_{0},\ l_{0}\in\mathbb{R}$, $b_{0}\phi_{1}^{*}+l_{0}\phi_{2}^{*}+\phi_{0,0}^{*}$ solves (2.28), and has boundary values $(\bar{p}(0),\ \bar{x}(0),\ {b}(0),\ {l}(0))^{\top}=(p_{0},\ \bar{x}_{0}^{N},\ b_{0},\ l_{0})^{\top}$. The uniqueness of solutions provided by subsection 2.2 indicates that $\phi_{b_{0},l_{0}}^{*}=b_{0}\phi_{1}^{*}+l_{0}\phi_{2}^{*}+\phi_{0,0}^{*}$. Therefore, by letting $t=T$, $\phi_{b_{0},l_{0}}^{*}(T)=b_{0}\phi_{1}^{*}(T)+l_{0}\phi_{2}^{*}(T)+\phi_{0,0}^{*}(T)$. By Assumption $(\boldsymbol{A_{3}})$, there exists a unique pair of constants $b_{0},\ l_{0}\in\mathbb{R}$ such that the initial values $(\bar{p}(0),\ \bar{x}(0),\ {b}(0),\ {l}(0))^{\top}=(p_{0},\ \bar{x}_{0}^{N},\ b_{0},\ l_{0})^{\top}$ gives a solution that satisfies (2.28) and (2.29). Thus, the solution for (2.28) and (2.29) exists and is uniquely determined. The proof is complete. ∎

### v2

@@ Appendix B Proofs of Auxiliary Propositions
Proof of section 3.
First, we need to confirm that (3.6) admits a unique classical solution on $[0,T]$. As stated before, we abbreviate $a_{i}$ and $B_{i}$ as $a$ and $B$, respectively. To prove existence, let function
@@
where $A(t)$ is a primitive function for $\frac{2a(t)}{c}$, so (3.7) gives a unique solution on $[0,T]$. The solution to (3.8) obviously exists and is unique, once we integrate both sides from $T$ to $t$. The proof is complete. ∎
Proof of section 3.
@@
We introduce the following notation: the Frobenius norm of any matrix $L$ is denoted by $\|L\|_{\mathrm{F}}$. It is clear that $\|M(t)\|_{\mathrm{F}}$ has a uniform bound for $t\in[0,T]$; we denote this bound by $M^{*}$, and denote $\|\bar{k}(t)\|_{\mathrm{F}}$ by $n(t)$. Multiply both sides of (3.9) with $2\bar{k}(t)^{\top}$, we obtain
@@
By ODE comparison theorems, $n(t)$ stays below a function $N(t)$ characterized by
@@
It directly follows that $N(t)$ has a uniform bound on $t\in[0,T]$, so $\bar{k}(t)$ stays finite on $[0,T]$. Therefore, the existence of the solution holds. Moreover, $M(t)$ is Lipschitz continuous w.r.t $t\in[0,T]$, so uniqueness naturally follows from the Picard-Lindelöf theorem. The proof is complete.
Proof of section 3.
We adopt the notation established in the proofs of section 3. Deducing in the same way as in the proof of section 3, we obtain the existence property. Recalling that $M(t)$ is Lipschitz continuous, we obtain the uniqueness property. ∎
Proof of section 3.
We adopt the notation established previously in section 3. By the definition of $\phi_{b_{0}}$, as shown in section 3 , we discover that for any $b_{0}\neq 0$, $\frac{\phi_{b_{0}}-\phi_{0}}{b_{0}}$ solves (3.11). Consequently, from the uniqueness property provided by section 3, we have
@@
Set $t=T$. From $B_{1}(T)\neq 0$, we conclude that there exists a unique $b_{0}\in\mathbb{R}$ that satisfies $\phi_{b_{0}}(T)=-\zeta\gamma$. Thus, there exists only one solution to (3.9) and (3.10). The proof is complete. ∎
Proof of section 4.
@@
To prove existence, we observe that $\|G(t)\|_{\mathrm{F}}$ has a uniform bound on $[0,T]$. Consequently, by working exactly as in section 3, we obtain the uniqueness property. Note that $G(t)$ is continuously differentiable, so the solution of the system is unique, from the Picard-Lindelöf theorem. ∎
Proof of section 4.
The existence holds by the same way as in section 3. Recalling from section 4 that $G(t)$ is continuously differentiable everywhere, we conclude that the solution is unique, by the Picard-Lindelöf theorem. ∎
Proof of section 4.
Observe that , for $\forall b_{0},\ l_{0}\in\mathbb{R}$, $b_{0}\phi_{1}^{*}+l_{0}\phi_{2}^{*}+\phi_{0,0}^{*}$ solves (4.11), and has boundary values $(\bar{p}(0),\ \bar{x}(0),\ {b}(0),\ {l}(0))^{\top}=(p_{0},\ \bar{x}_{0}^{N},\ b_{0},\ l_{0})^{\top}$. Then, uniqueness property in section 4 yields that $\phi_{b_{0},l_{0}}^{*}=b_{0}\phi_{1}^{*}+l_{0}\phi_{2}^{*}+\phi_{0,0}^{*}$. Therefore, by setting $t=T$, we get $\phi_{b_{0},l_{0}}^{*}(T)=b_{0}\phi_{1}^{*}(T)+l_{0}\phi_{2}^{*}(T)+\phi_{0,0}^{*}(T)$. By Assumption $(\boldsymbol{A_{3}})$, there exists a unique pair of constants $b_{0},\ l_{0}\in\mathbb{R}$ such that the initial values $(\bar{p}(0),\ \bar{x}(0),\ {b}(0),\ {l}(0))^{\top}=(p_{0},\ \bar{x}_{0}^{N},\ b_{0},\ l_{0})^{\top}$ gives a solution that satisfies (4.11) and (4.12). Thus, the solution for (4.11) and (4.12) exists and is unique. The proof is complete. ∎